OpenStudy (anonymous):

Tammy is at the dentist's office waiting on her appointment. She notices that the 6-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle. Part 1: How many radians does the minute hand move from 1:20 to 1:55? (Hint: Find the number of degrees per minute first.) Part 2: How far does the tip of the minute hand travel during that time? Part 3: How many radians on the unit circle would the minute hand travel from 0° if it were to move 5π inches? Part 4: What is the coordinate point associated with this radian measure? show all of your work

OpenStudy (mathmale):

Let's paraphrase this Part 1: a) What is the "central angle" when the minute had traces one complete circle? b) What is the central angle when the tip of the minute hand moves from time 1:20 to time 1:55? c) Use the 'arc length' formula, s = r * theta, to answer the question posed in Part !.

OpenStudy (anonymous):

ok. now how do i find the number of degrees per minute

OpenStudy (phi):

Part 1: How many radians does the minute hand move from 1:20 to 1:55? can you answer how many minutes (of time) the hand moved between 1:20 and 1:55 ?

OpenStudy (anonymous):

35 minutes

OpenStudy (phi):

a full circle is 60 minutes. what fraction is 35 minutes of the full circle ?

OpenStudy (anonymous):

hmmm i know its over half

OpenStudy (phi):

fraction will be part over whole

OpenStudy (anonymous):

so probably 5/8? or is it a larger fraction

OpenStudy (phi):

35 minutes out of 60 minutes, written as a fraction

OpenStudy (anonymous):

i think its 7/12

OpenStudy (phi):

it's the same idea as if you had 3 spoons, and you lose 2 of them. you lost ⅔ of your spoons. part/whole or 2/ 3 here 35/60

OpenStudy (phi):

yes, 7/12 next idea: there are 2 pi radians in a full circle. you have 7/12 of a full circle how many radians do you have?

OpenStudy (phi):

or, if you like to first use degrees. There are 360 degrees in a full circle and you have 7/12 of a full circle. how many degrees? (then to get radians multiply by 2 pi/180)

OpenStudy (phi):

***oops: (then to get radians multiply by pi/180)

OpenStudy (anonymous):

210 degrees

OpenStudy (phi):

if the fraction were ½, it's more obvious. full circle is 2 pi radians ½ of that is ½ * 2pi = pi if you have ¾ of a circle, you have ¾ * 2 pi = 3 pi/2 yes 210º

OpenStudy (anonymous):

and then i multiply that by pi/180?

OpenStudy (phi):

yes

OpenStudy (phi):

you should get 7/12 * 2 pi = 7pi/6 radians

OpenStudy (anonymous):

ok... i got 7pi / 6

OpenStudy (anonymous):

ok good

OpenStudy (phi):

btw, their hint meant: 1 minute is 1/60 of a full circle. in degrees 1/60 of 360º= 6º per minute then multiply 35 minutes times 6º/minute to get 210º that works too (obviously)

OpenStudy (anonymous):

ooh ok i didnt see that lol

OpenStudy (phi):

Part 2: How far does the tip of the minute hand travel during that time? the minute hand's tip moved along the circumference of the circle. It went 7/12 of "all the way around" so one way to do the problem is figure out the circumference (using the very famous C= 2 pi r formula), then multiply by 7/12 or , because we know the angle in radians, we use the simple formula: arc= r theta, where theta is the angle in radians

OpenStudy (anonymous):

ok lemme finish filling out part 1 real quick

OpenStudy (anonymous):

ok done now lemme look at pt. 2

OpenStudy (phi):

what's a lemme?

OpenStudy (anonymous):

"let me" but its shorthand

OpenStudy (phi):

oh

OpenStudy (anonymous):

ok in the C=2pi r... our r is our radius correct

OpenStudy (phi):

yes

OpenStudy (anonymous):

would our radius be 6 inces because it says she notices the minute hand is 6 inches long

OpenStudy (phi):

the radius of a circle is the distance from the center to the circumference. the minute hand starts at the center and goes to the circumference. its length is the radius

OpenStudy (anonymous):

ok thats what i thought. so its C= 2(pi)(6)

OpenStudy (phi):

yes, which simplifies to 12 pi

OpenStudy (anonymous):

then i multiply that by 7/12

OpenStudy (phi):

yes, because that is the fraction of the whole circumference the hand moved

OpenStudy (anonymous):

i got 7pi

OpenStudy (phi):

yes. based on the question, they may want you to use the formula \[ s = r \ \theta \] where r is 6, and theta is the angle in radians (from Part I) s is the length of the arc in inches. try that way.

OpenStudy (anonymous):

our angle for theta would be 35?

OpenStudy (phi):

no, theta is the answer to part I.

OpenStudy (anonymous):

ok so our 7pi/6? or the 210

OpenStudy (phi):

remember, the minute hand moved from 1:20 to 1:55 which is 35 minutes which is 35/60 of 360 degrees , which is 210º, which is 7pi/6 radians

OpenStudy (phi):

the formula only works for theta in *radians*

OpenStudy (anonymous):

oohhhh ok gotcha

OpenStudy (anonymous):

ummm i got 7pi? is that right

OpenStudy (phi):

does that match what you got doing it the other way ?

OpenStudy (anonymous):

what do you mean by the other way?

OpenStudy (phi):

scroll up

OpenStudy (anonymous):

yeah i did it in s=r(theta)

OpenStudy (phi):

you also did C= 2 pi r = 2 * pi*6= 12 pi then 7/12 * 12 pi = 7 pi

OpenStudy (anonymous):

yes and i got the same answer for both

OpenStudy (phi):

The first way is more common sense. the second way is faster if (1) you remember the formula (2) have theta in radians

OpenStudy (anonymous):

ok id have to say it did seem quicker. alot less to write and was simpler

OpenStudy (phi):

Part 3: How many radians on the unit circle would the minute hand travel from 0° if it were to move 5π inches?

OpenStudy (anonymous):

almost done filling out pt. 2

OpenStudy (anonymous):

done

OpenStudy (phi):

For part 3, I would use s=r θ they give you s (the distance the minute hand moved), and r= 6 so you can find theta (it will be in radians)

OpenStudy (anonymous):

would our s be 35?

OpenStudy (phi):

35 is minutes. we are now looking at a circle with a radius of 6 inches and talking about moving so many inches around it (i.e. along its circumference) they told us how far we moved.

OpenStudy (anonymous):

which was 7/12 or 210 degrees?

OpenStudy (phi):

yes, but part 3 has moved to a different question. Part 3: How many radians on the unit circle would the minute hand travel from 0° if it were to move 5π inches? in other words, the minute hand moved 5π inches

OpenStudy (anonymous):

soo multiply 5pi to see how many inche it moved?

OpenStudy (phi):

the key thing is it moved *inches*. we could get a ruler and try to measure along the circle about 15.7 inches, to see where it ended up but they use the goofy 5 pi (instead of a number) so the numbers will work out nicer. So keep the distance as 5 pi (rather than 15.70796327....)

OpenStudy (phi):

The idea is the *length of the arc* is measured in inches we use s=r θ

OpenStudy (phi):

s=r θ they give you s (the distance the minute hand moved), and r= 6 so you can find theta (it will be in radians)

OpenStudy (anonymous):

ok i had to step away for a sec lemme catch up real quick

OpenStudy (anonymous):

ok so s would be 5 pi = 6(7pi/6)

OpenStudy (phi):

almost. they gave you s and r, but we don't know theta for the s they gave us. Remember: this is a different problem

OpenStudy (anonymous):

yep thats right i keep forgetting its a different problem so how would we find our theta

OpenStudy (phi):

first write down the correct equation

OpenStudy (anonymous):

ok 5pi = 6(theta)

OpenStudy (phi):

any ideas how to "solve" for theta?

OpenStudy (phi):

use these two ideas: 1) 6/6 is 1 2) if we divide one side of the equation by 6, we also divide the other side by 6

OpenStudy (anonymous):

i cant find anything about it in my notes but i am missing 2 pages... i need to keep better notes lol

OpenStudy (anonymous):

ok i get that

OpenStudy (anonymous):

so what we do to one side we do to the other

OpenStudy (phi):

yes. and in this case we want to divide the right side by 6 because it makes the 6/6 = 1

OpenStudy (anonymous):

so the didvide 5(pi) by 6?

OpenStudy (phi):

yes. not hard write 5pi/6 simple! you get 5 pi/6 radians = theta (you should put in the units so we don't get confused about inches, degrees or radians) remember this formula s= r theta assumes theta is in *radians*

OpenStudy (anonymous):

so 5pi divided by 6 and ur answer is theta or theta is 5pi/6 the equation

OpenStudy (phi):

can you be more clear?

OpenStudy (anonymous):

i know it confused me

OpenStudy (phi):

From the beginning (of Part 3) Part 3: How many radians on the unit circle would the minute hand travel from 0° if it were to move 5π inches? that says the minute hand moved 5 pi inches (along the arc) they want to know what angle that is (but in radians not degrees) we use s = r theta where s is the arc in inches, r is the radius in inches, theta is the angle in radians

OpenStudy (anonymous):

ok theta be the sum of the nequation 5pi/6 or would theta be the equation 5pi over 6

OpenStudy (phi):

you found 5 pi/6 radians = theta in other words, the angle theta is 5 pi/ 6 radians that is the answer.

OpenStudy (phi):

How many radians on the unit circle would the minute hand travel? 5 pi/ 6 radians

OpenStudy (anonymous):

ok so it would be 5pi = 6 (5pi/6)

OpenStudy (phi):

that is a true equation (simplify the right side to get 5 pi) but the idea is 1) we want to find the angle theta 2) we found theta= 5 pi / 6 radians we found what the angle is. use its value as the answer to the question How many radians on the unit circle would the minute hand travel?

OpenStudy (anonymous):

ok so part 3 answer is 5pi/6 radians

OpenStudy (phi):

yes

OpenStudy (anonymous):

ok sorry im at work right now while doing classes lol. but im back

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

one last part